OpenStudy (anonymous):

The graph of which quadratic equation has a vertex of (–2, 1)? Answer y = (x + 1)^2 – 2 y = (x + 2)^2 + 1 y = (x – 2)^2 + 1 y = (x – 1)^2 – 2 The graph of which quadratic equation has a vertex of (–2, 1)? Answer y = (x + 1)^2 – 2 y = (x + 2)^2 + 1 y = (x – 2)^2 + 1 y = (x – 1)^2 – 2 @Mathematics

OpenStudy (anonymous):

y = (x-2)^2 + 1

OpenStudy (anonymous):

(h, k) --> (-2, 1) y = (x-h)^2 + k y = (x - 2)^2 + 1